OpenStudy (anonymous):

why is it that when 'using intercepts and symmetry to sketch a graph' does the equation x-y^2=1 get the y replaced w/ a negative [looking like this -> x-(-y)^2=1 ] wind up being simplified to x-y^2=1 instead of being simplified to x+y^2=1 ?

OpenStudy (dan815):

what thats messed up forget all about that stuff

OpenStudy (anonymous):

aw, well :( unfortunately I can not because it is a fundamental question for me to understand the section but thanks

OpenStudy (jhannybean):

\[\large x-(-y)^2 = 1 \] solve the (-y)^2 first. \[\large (-y^2) = (-y)(-y) = y^2\]now replace (-y)^2 with y^2\[\large x-y^2 = 1\]

OpenStudy (jhannybean):

The y doesn't get replaced by a negative, you just have to remember to solve for your (-y)^2 first before simplifying x-(-y)^2 into x+y^2.

OpenStudy (anonymous):

thank you, wowww I cant believe I didn't catch that

OpenStudy (dan815):

i said forget all this stuff because i dont like how ur thinking about equations

OpenStudy (dan815):

we dont think about all this stuff, when ur just building a vertex equations, u just wanna kinda build ur equation knowing where the Vertex will be

OpenStudy (dan815):

dont be restrained to these rules, learn to adapt then u will find math easier as u get better at adapting to different methods

OpenStudy (anonymous):

okidiki, thank you foor that insight

OpenStudy (anonymous):

okidoki, lol